OpenStudy (anonymous):

A barrel contains 149 gallons of water and is being drained at a constant rate of 15 gallons per hour. Write the equation that modles the number of gallons,g, after t hours.

OpenStudy (anonymous):

how to get the equation\?

OpenStudy (anonymous):

why g?

OpenStudy (anonymous):

for gallons?

OpenStudy (unklerhaukus):

initial condition \[g(0)=g_0=149\left[ {\text{gallons of}\text{ H}_2\text O} \right]\] rate of change \[\frac{\text d g(t)}{\text dt}=-15 \left[\frac {\text{gallons}\text{ H}_2\text{O}}{\text {hour}}\right]\]\[\downarrow\]\[{\text d g(t)}=-15 \cdot\text dt\]\[{\int\text d g(t)}=\int-15 \cdot\text dt\]\[g(t)=-15\int\text dt\]

OpenStudy (unklerhaukus):

you'll be able to find the constant \(c\) of integration by substituting in \[g(0) \] at \[t=0\]

OpenStudy (unklerhaukus):

@unkabogable !